# Model-Based Transfer Learning for Contextual Reinforcement Learning

### v1

@@ A.1.1 Notation
$\mathbf{x}$ Vector of source tasks ($\mathbf{x}=[x_{1},...,x_{k}]$ where $x_{i}\in X$ for $1\leq i\leq k$)
$c\mathcal{M}(x)$ Contextual MDP parameterized by $x$
$J(x)$ Performance of task $c\mathcal{M}(x)$
$\Delta\hat{J}(x,x^{\prime})$ Generalization gap (source: $x$, target: $x^{\prime}$)
$U(x^{\prime};x)$ Transferred performance (source: $x$ (or $\mathbf{x}$), target: $x^{\prime}$)
$V(x)$ Expected generalized performance of source model $x$ evaluated on all $x^{\prime}\in X$
Figure 8 helps understand the discrepancy between the observed generalized performance and the predicted one. Figure 9 illustrates how to calculate the marginal improvement of expected generalized performance ($\hat{V}(x;x_{1},...,x_{k-1})-V(x_{1},...,x_{k-1})$).
@@
Figure 8: Illustration of the discrepancy between observed ($U$) and predicted ($\hat{U}$) generalized performance after training on source task $x_{1}$ and attempting zero-shot transfer to $x^{\prime}$.
@@
Figure 9: Step for choosing $x_{2}$ that maximizes the estimated marginal improvement ($\hat{V}(x;x_{1})-V(x_{1})$). $\hat{V}(x;x_{1})$ corresponds to the red area under the red line and $V(x_{1})$ as the area under $U(x^{\prime};x_{1})$.
A.1.2 Proof of Theorem 2
@@
A.2 Experiment details
A.2.1 Details about Gaussian process (GP)
In our study, we conducted hyperparameter tuning experiments for Gaussian Process (GP) regression to optimize its performance. Specifically, we varied the noise standard deviation over the set $\{0.001,0.01,0.1,1\}$ and the number of restarts for the optimizer over the set $\{5,10,15\}$. We explored several kernel configurations, including the combination of a constant kernel (C) and a radial basis function kernel (RBF) with a length scale ranging from 0.01 to 100, as well as more complex kernels incorporating a white noise kernel and an exponential sine squared kernel with varying length scales and periodicities. Each GP model was instantiated with the specified kernel, and the noise level was set to the square of the noise standard deviation. This systematic exploration aimed to identify the optimal set of hyperparameters that could enhance the model’s predictive accuracy.
A.2.2 Results of table with standard deviation
Table 4: Comparative performance of different methods on context-variant traffic and benchmark tasks
Benchmark Baselines Model-Based Transfer Learning Oracle
Task Context Variation Random Exhaustive Multitask MBTL-PS MBTL-ES MBTL-GP Oracle Transfer
Number of Trained Models $k$ $N$ $1$ $k$ $K$ $k$ $N$
@@ A.2.2 Results of table with standard deviation
* Note: The bold values indicate the maximum value for each task excluding the oracle. Standard deviation across multiple runs in the parenthesis.
A.2.3 Details about traffic signal control
@@ Results
  • Inflow: Performance improves as the number of transfer steps increases, with MBTL-GP strategy consistently achieving the highest scores, demonstrating their effectiveness in adapting to changes in inflow conditions.
@@
  • Road Length: There is a general upward trend in performance for all strategies, particularly for MBTL-GP, indicating robustness in adapting to different road lengths.
This data suggests that MBTL-GP and Oracle are particularly effective across varying conditions, maintaining higher levels of performance adaptability.
@@
A.2.4 Details about eco-driving Lagrangian control
Given the significant portion of greenhouse gas emissions in the United States coming from the transportation sector [11], eco-driving behaviors are critical for climate change mitigation. Deep reinforcement learning-based eco-driving strategies have been developed [13, 38, 17] but still have some issues of difficulties in generalization. We also extend to various intersection configurations with different traffic inflow rates, penetration rates of eco-driving systems, and durations of green phases at static traffic signals to optimize vehicle behaviors for reduced emissions.
Figure 13 illustrates the traffic road network used in eco-driving Lagrangian control task. The road network is depicted with traffic flowing vertically and horizontally crossing the static phase traffic signal. There are both guided and default vehicles in the system. The state space includes the speed and position of the ego vehicle, the leading vehicle, and the following vehicles, supplemented by the current traffic signal phase and relevant context features, including lane length and green phase durations. The action space specifically focuses on the ego vehicle’s acceleration control. The reward mechanism is designed to optimize the driving strategy by balancing the average speed of the vehicles against penalties for emissions, thereby promoting eco-friendly driving behaviors within the traffic system.
@@ A.2.4 Details about eco-driving Lagrangian control
Figure 14: Examples of transferability heatmap for eco-driving Lagrangian control.
@@ Results
Figure 15 illustrates the normalized generalized performance across variants of eco-driving control tasks, specifically looking at variations in green phase time, inflow, and penetration rate. The graphs depict performance enhancement over transfer steps for different strategies. Notably, MBTL-GP consistently demonstrates superior performance across all variations, indicating robust adaptability to changing task parameters.
Figure 15: Comparison of normalized generalized performance of all target tasks: Eco-driving Lagrangian control.
A.2.5 Details about advisory autonomy task
Advisory autonomy involves a real-time speed advisory system that enables human drivers to emulate the system-level performance of autonomous vehicles in mixed autonomy systems [34, 15]. Instead of direct and instantaneous control, human drivers receive periodic guidance, which varies based on road type and guidance strategy. Here, we consider the different frequencies of this periodic guidance as contextual MDPs since the zero-order hold action affects the transition function.
@@
Figure 18 illustrates the comparison of normalized generalized performance for advisory autonomy tasks, specifically acceleration and speed guidance in a ring and acceleration guidance on a ramp. The graphs demonstrate that MBTL-GP consistently exhibits higher performance across all tasks. Particularly, acceleration guidance in both ring and ramp scenarios shows significant performance improvements over transfer steps, with MBTL-GP closely matching in some instances.
@@
A.2.6 Details of control tasks
@@
A.2.7 Results of control tasks
Figure 19 compares our proposed methods with baselines and oracle method in different variations of context in classic control tasks.
Figure 19: Comparison on control benchmarks: Cartpole, Pendulum, BipedalWalker, and Halfcheetah.
A.2.8 Details about Cartpole
@@ Results
In the mass of cart variation, the performance generally increases with transfer steps, with MBTL-GP strategies achieving the highest scores, indicating robust adaptability to changes in cart mass. Similar trends are observed with length variation and mass of pole variation. MBTL-GP shows close to oracle performance.
@@
A.2.9 Details about Pendulum
@@
Figure 23 shows a comparison of normalized generalized performance for the Pendulum task across different strategies when varying the timestep, length of the pendulum, and mass of the pendulum. For the length of the pendulum variation and mass of the pendulum one, MBTL-GP strategies demonstrate the highest scores, suggesting robust adaptability to changes in pendulum dynamics. MBTL-GP shows performance close to that of the Oracle across all variations, indicating its effectiveness in handling dynamic changes in system parameters.
@@
A.2.10 Details about BipedalWalker
@@
Figure 25 shows the comparison of normalized generalized performance for all variations within the BipedalWalker task. There is no huge difference in performance for all three cases, but if we look into the tabualr results in Table 2, MBTL-GP shows the highest performance across varying conditions, indicating their robustness in adapting to changes in physical parameters of the model. This suggests that these strategies are more effective in handling the complexities introduced by different frictions, gravities, and scales compared to other baselines.
@@
A.2.11 Details about HalfCheetah
@@
Figure 27 presents a comparison of normalized generalized performance across various strategies for the HalfCheetah task with respect to the varied physical properties. The results indicate that the MBTL-GP generally outperforms others, particularly in managing variations in gravity and stiffness, suggesting the superior adaptability of these models to physical changes in the task environment. The trends across different parameters confirm the critical impact of task-specific dynamics on the effectiveness of the tested strategies.

### v2

@@ A.1 Notation
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}}$ Trained policy from source task ($x\in X$)
@@
$\mathcal{M}_{x}$ Contextual MDP parameterized by $x$
$J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x)$ Performance of task $\mathcal{M}_{x}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J(\pi_{x},x^{\prime})}$ Generalization performance (source: $x$ (or $\mathbf{x}$), target: $x^{\prime}$)
$\Delta{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},}x^{\prime})$ Generalization gap (source: $x$, target: $x^{\prime}$)
$V(x^{\prime};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}})$ Expected generalization performance of source model $x$ evaluated on all $x^{\prime}\in X$
Figure 9 helps understand the discrepancy between the observed generalized performance and the predicted one. Figure 10 illustrates how to calculate the marginal improvement of expected generalized performance ($\hat{V}(x;x_{1:k-1})-V(x;x_{1:k-1})$).
@@
Figure 9: Illustration of the discrepancy between observed ($J$) and predicted ($\hat{J}$) generalized performance after training on source task $x_{1}$ and attempting zero-shot transfer to $x^{\prime}$.
@@
Figure 10: Step for choosing $x_{2}$ that maximizes the estimated marginal improvement ($\hat{V}(x;\pi_{1})-V(x_{1})$). $\hat{V}(x;\pi_{1})$ corresponds to the red area under the red line and $V(x_{1})$ as the area under $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1},x^{\prime}})$.
A.2 Model-Based Transfer Learning (MBTL) Algorithm
[…]
A.3 Theoretical analysis
A.3.1 Proof of Theorem 2
@@
A.4 Experiment details
A.4.1 Details about Gaussian process (GP)
In our study, we conducted hyperparameter tuning experiments for Gaussian Process (GP) regression to optimize its performance. Specifically, we varied the noise standard deviation over the set $\{0.001,0.01,0.1,1\}$ and the number of restarts for the optimizer over the set $\{5,6,7,8,9,10,11,12,13,14,15\}$ with several kernel configurations. We set noise standard deviation as $0.1$, the number of restarts for the optimizer over the set as $9$, and the kernel as the combination of a constant kernel (C) with the bound from 0.001 to 1000 and a radial basis function kernel (RBF) with a length scale ranging from 0.01 to 100. The noise level was set to the square of the noise standard deviation ($0.1^{2}$).
A.4.2 Results of table with standard deviation
Table 4: Comparative performance of different methods on context-variant traffic and benchmark CMDPs
Benchmark (CMDP) Baselines Sequential Baselines MBTL Oracle
Domain Context Variation Independent Multi-task Random Equidistant Greedy Ours Sequential
Number of Trained Models $N$ $1$ $k$ $K$ $k$ $k$ $N$
@@ A.4.2 Results of table with standard deviation
* Note: Bold values represent the highest value(s) within the statistically significant range for each task, excluding the oracle. Standard deviation across multiple runs in the parenthesis.
A.4.3 Detailed sample complexity comparison results
Table 5 presents a comparison of sample complexity required for MBTL to perform as good as the best generalization performance of baselines (Independent training and multi-task training) across various tasks in the CMDP. Each row lists a different domain, the specific context variation applied (e.g., changes in physical properties or environmental parameters), and two key values: $k^{*}$ and $N$, where $k^{*}$ represents the number of models required by MBTL to reach a performance level comparable to the baseline. This value is shown as a range (e.g., $[3,5,3]$), indicating results from three random seeds. $N$ represents the total number of contexts. The value $\frac{N}{k^{*}}$ helps illustrate the sample efficiency of MBTL.
Table 5: Sample complexity comparison to baseline performance on CMDP tasks
Task Context Variation $k^{*}$ $N$
Pendulum Length [3, 5, 3] 100
Pendulum Mass [3, 3, 3] 100
Pendulum Timestep [6, 5, 6] 100
Cartpole Mass of Cart [9, 13, 6] 100
Cartpole Length of Pole [4, 4, 3] 100
Cartpole Mass of Pole [3, 3, 3] 100
BipedalWalker Gravity [2, 2, 3] 100
BipedalWalker Friction [2, 2, 2] 100
BipedalWalker Scale [3, 3, 1] 100
HalfCheetah Gravity [2, 2, 2] 100
HalfCheetah Friction [1, 2, 2] 100
HalfCheetah Stiffness [1, 2, 2] 100
AA-Ring-Acc Hold Duration [3, 8, 2] 40
AA-Ring-Vel Hold Duration [2, 1, 1] 40
AA-Ramp-Acc Hold Duration [1, 2, 2] 40
AA-Ramp-Vel Hold Duration [2, 2, 2] 40
Traffic Signal Road Length [8, 5, 6] 50
Traffic Signal Inflow [2, 2, 2] 50
Traffic Signal Speed Limit [4, 4, 5] 50
Eco-Driving Penetration Rate [2, 2, 1] 50
Eco-Driving Inflow [2, 1, 1] 50
Eco-Driving Green Phase [2, 2, 2] 50
A.4.4 Details about traffic signal control
@@ Results
  • Inflow: Performance improves as the number of transfer steps increases, with MBTL strategy consistently achieving the highest scores, demonstrating their effectiveness in adapting to changes in inflow conditions.
@@
  • Road Length: There is a general upward trend in performance for all strategies, particularly for MBTL, indicating robustness in adapting to different road lengths.
This data suggests that MBTL and Oracle are particularly effective across varying conditions, maintaining higher levels of performance adaptability.
@@
A.4.5 Details about eco-driving control
Given the significant portion of greenhouse gas emissions in the United States coming from the transportation sector [12], eco-driving behaviors are critical for climate change mitigation. Deep reinforcement learning-based eco-driving strategies have been developed [14, 45, 18, 20] but still have some issues of difficulties in generalization. We also extend to various intersection configurations with different traffic inflow rates, penetration rates of eco-driving systems, and durations of green phases at static traffic signals to optimize vehicle behaviors for reduced emissions.
Figure 14 illustrates the traffic road network used in the eco-driving control task. The road network is depicted as traffic flowing vertically and horizontally, crossing the static phase traffic signal. There are both guided and default vehicles in the system. The state space includes the speed and position of the ego vehicle, the leading vehicle, and the following vehicles, supplemented by the current traffic signal phase and relevant context features, including lane length and green phase durations. The action space specifically focuses on the ego vehicle’s acceleration control. The reward mechanism is designed to optimize the driving strategy by balancing the average speed of the vehicles against penalties for emissions, thereby promoting eco-friendly driving behaviors within the traffic system.
@@ A.4.5 Details about eco-driving control
Figure 15: Examples of transferability heatmap for eco-driving control.
@@ Results
Figure 16 illustrates the normalized generalized performance across variants of eco-driving control tasks, specifically looking at variations in green phase time, inflow, and penetration rate. The graphs depict performance enhancement over transfer steps for different strategies. Notably, MBTL consistently demonstrates superior performance across all variations, indicating robust adaptability to changing task parameters.
Figure 16: Comparison of normalized generalized performance of all target tasks: Eco-driving control.
A.4.6 Details about advisory autonomy task
Advisory autonomy involves a real-time speed advisory system that enables human drivers to emulate the system-level performance of autonomous vehicles in mixed autonomy systems [39, 8, 16]. Instead of direct and instantaneous control, human drivers receive periodic guidance, which varies based on road type and guidance strategy. Here, we consider the different frequencies of this periodic guidance as contextual MDPs since the zero-order hold action affects the transition function.
@@
Figure 19 illustrates the comparison of normalized generalized performance for advisory autonomy tasks, specifically acceleration and speed guidance in a ring and acceleration guidance on a ramp. The graphs demonstrate that MBTL consistently exhibits higher performance across all tasks. Particularly, acceleration guidance in both ring and ramp scenarios shows significant performance improvements over transfer steps, with MBTL closely matching in some instances.
@@
A.4.7 Details of control tasks
@@
A.4.8 Details about Cartpole
@@ Results
In the mass of cart variation, the performance generally increases with transfer steps, with MBTL strategies achieving the highest scores, indicating robust adaptability to changes in cart mass. Similar trends are observed with length variation and mass of pole variation. MBTL shows close to oracle performance.
@@
A.4.9 Details about Pendulum
@@
Figure 23 shows a comparison of normalized generalized performance for the Pendulum task across different strategies when varying the timestep, length of the pendulum, and mass of the pendulum. For the length of the pendulum variation and mass of the pendulum one, MBTL strategies demonstrate the highest scores, suggesting robust adaptability to changes in pendulum dynamics. MBTL shows performance close to that of the Oracle across all variations, indicating its effectiveness in handling dynamic changes in system parameters.
@@
A.4.10 Details about BipedalWalker
@@
Figure 25 shows the comparison of normalized generalized performance for all variations within the BipedalWalker task. There is no huge difference in performance for all three cases, but if we look into the tabualr results in Table 2, MBTL shows the highest performance across varying conditions, indicating their robustness in adapting to changes in physical parameters of the model. This suggests that these strategies are more effective in handling the complexities introduced by different frictions, gravities, and scales compared to other baselines.
@@
A.4.11 Details about HalfCheetah
@@
Figure 27 presents a comparison of normalized generalized performance across various strategies for the HalfCheetah task with respect to the varied physical properties. The results indicate that the MBTL generally outperforms others, particularly in managing variations in gravity and stiffness, suggesting the superior adaptability of these models to physical changes in the task environment. The trends across different parameters confirm the critical impact of task-specific dynamics on the effectiveness of the tested strategies.